Operations Report – March 05th

Crew 223 Operations Report 05/03/2020

SOL: 4

Name of person filing report: Luc FORTIN

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Wednesday generator:
Off at 6:00am
On at 9:15pm SOC was 84%

Thursday generator:
Off at 4:30am
On at 9pm SOC was XX%

Diesel Reading – 70%

Station Propane Reading – 82%

Ethanol Free Gasoline – 7 gallons

Water (loft tank): 47 gallons

Water Meter: 149709.09 units

Water (static tank) – 319 gallons

Static to Loft Pump used – yes

Water in GreenHab – 166.4 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: yes

Perseverance (former Demos) rover used: no
Hours: 149.5
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Spirit rover: yes
Hours: 133.7
Beginning charge: 100%
Ending charge: 86%
Currently charging: yes

Opportunity rover used: moved from campus to the shop to repair the brakes.
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Curiosity rover used: no
Hours: 139.3
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: yes (Honda, 1, 2 and 3)
Reason for use: 1 & 2 used during the EVA. We had trouble starting the
Honda, and finally thought the battery was flat. But during the engineer
EVA, Honda started without any problem (using the starter). So maybe it
was due to the cold weather… Crew members going in EVA tomorrow
morning will try to start it. If non working, we will probably take the
battery in the RAM to charge it. All ATVs were run during the engineer
EVA for battery check.
Oil Added? no
ATV Fuel Used: ~5%
Hours the ATVs were Used today: 20 minutes
Notes on ATVs: Nothing to report

HabCar used and why, where? Nothing to report

CrewCar used and why, where? The CrewCar is in Grand Junction.

General notes and comments: Nothing to report

Summary of internet: All ok.

Summary of suits and radios: Suits 4, 8, 9, 10 were used during the EVA.
Suit 7 is now fixed and perflectly working !

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: The GreenHab officier switched the
working and non working neons of the halogen lamp in order to have more
light on tomatoes than on aromatics.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: We successfully fixed suit 7 ! What we found
is a problem on the suit female plug : in a certain position, one of the
cables was in contact with one of the 2 free pins. And thus connecting
the electrical circuit to the ground. Our solution was to isolate those
2 pins with an heat-shrinkable sleeve. Cf the 2 pictures attached. We
did the same thing on suits 8, 9, 10 to prevent them from this.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

Sol Summary – March 05th

Crew 223 Sol Summary Report 05-03-2020

Sol: 4 Summary Title: Fisrt routine day

Author’s name: Aurélien Mure – Crew Commander

Mission Status: Crew 223 Routine has begun

Sol Activity Summary:

The crew woke up at 6:50am and every. The sport session began at 7:00am.
It seems that the whole team is feeling more confident with the sport
session and is improving the performances. It is good to see.
Marion was the EVA leader for the first time, she was a little bit
stressed out but she managed the EVA prep without any problem. The ATV
Honda refused to start at the beginning of the EVA. Whe chose to pick
the ATV 3 instead. We explored Candor Chasma through Stream Bed Connector.
Marion’s spacesuit was not really well set so that she lost her
earphone. We found a lot of garages there and we tried to pick it up as
much as possible.
Once arrived at the Hab we tried to start the ATV Honda manually but
without success.
Good news for the end of the morning : The Spacesuit 7 is fixed and we
have found a possible weakness on the spacesuit design that we have
fixed on the three other suits. I worked with Luc 3 hours on it, we hope
that everything will be all right tomorrow.
Florian worked on his solar imaging, Blandine and Marion on their app
displaying the schedule. We continued with another cohesion activity in
Blandine’s program.

Look Ahead Plan:

The suits are going to be tested tomorrow, we hope to go further with
our solar oven.

Anomalies in work: None.

Weather: Sunny and warm all day (-6°C at 7:00am, about 9°C at 11:00am,
6:00pm) 0m/s of wind all day

Crew Physical Status: All crew members safe and sound.

EVA: LOAC and MegaAres batteries changed Candor Chasma exploration

Reports to be filed: Journalist Report, Operations report, EVA Request,
Green Hab report, Astronomy report, EVA report

Support Requested: None

Astronomy Report – March 05th

Crew 223 Astronomy Report Mar 05 2020

Name: Florian Delpech
Crew: 223
Date: 3/5/2020

MDRS ROBOTIC OBSERVATORY
Robotic Telescope Requested (choose one) MDRS-14

Objects to be Imaged this Evening: NGC 1961 ; NGC 4490 ; NGC 4450 ; NGC
3368 ; NGC 3115 ; NGC 2403 ; NGC 4618 ; NGC 4559 ; NGC 3521 ; NGC 4569 ;
NGC 4725 ; NGC 3675. This is part of the research project aiming at
observing about 30 galaxies per week to try to find a supernovae.

Images submitted with this report: N/A

Problems Encountered: N/A

MUSK OBSERVATORY

Solar Features Observed: Granules

Images submitted with this report: A picture of the Sun

Problems Encountered: N/A

GreenHab Report – March 05th

Crew 223 GreenHab Report – 05-03-2020 GreenHab Officer: Valentin
BELLEMIN-LAPONNAZ

Environmental control: Ambient with Door open

Average temperature: 23.2°C
Low temperature: 20°C High temperature: 26°C

Hours of supplemental light: 4 hours (from 9P.M to 1 a.m.)

Daily water usage for crops: 15.8 gallons

Daily water usage for research and/or other purposes: 0 gallon

Water in Blue Tank— 166.4 gallons

Time(s) of watering for crops: 3 (1 only for tomatoes and radish)

Changes to crops: I planted green spinach to replace the ones that I
harvested two days ago.

Narrative: This morning, I watered all the plants as usual. I planted
spinach in the free pot under the strawberries. I also did the cleaning
and marked well all the plants with new tags. After lunch, I carried out
the experiment Music For Plants. The radish which were transplanted
yesterday didn’t grow well. I had to find another solution. I decided to
transplant other advanced plants of radish. I hope those ones will
recover from transplantation. For the dinner, I harvested basil to make
a soup and lemon balm for the dessert.

Harvest: 10g Basil / 5g Lemon Balm

Support/supplies needed: NA

Journalist Report – March 05th

Crew 223 Journalist Report 05Mar2020
Author: Clément Plagne, Crew Journalist
Sol 4
Title: A matter of time
It’s sol 4 already! We’re just about a third of the way through our stay
in the station, which is honestly really hard to imagine. We’ve
experienced a lot already. The team has already changed from how it was
when we started, and we’ve grown as a unit more and more each day. Our
bodies are a bit leaner, bruised for a few of them, but feel stronger
and healthier – the all-American fast-food based diet we had before
arriving here is long gone. They’re also a whole lot dirtier. While no
one’s complained about smells, the days inside an enclosed, dusty
environment have taken a toll on us, and we’re all overdue for a shower.

Yeah, we sure have done a lot already. It’s hopeless to think we’ll
explore everywhere around us and do all the science we want, but we’ve
done a fair bit. Still, it feels like all those days have gone by like a
flash. Time goes by differently in the station. Saying “It’s sol 4!”
sounds fine to all of us, but “It’s Thursday!” just feels weird. Very
easily, you get into a rhythm. 7am wake up time, timed sports session,
breakfast, EVA Prep, check, check, check, 5 minutes of airlock
depressurisation, and on with the procedures. Everyone knows what
they’re doing every single minute of their morning. It’s doing work of
course, but also a lot of waiting for lunch. Lunch, as well as the
afternoon, is a lot freer. Where EVA mornings have us all focused on one
task, afternoons are for our own work. You get to take time to think.

When you’re in here, the rest of the world just stops to exist, and so
does everything with it. The days become simpler. No more social media,
no more advertisements on walks outside, no more commute to work.
Without distractions, you finally get to take time. Time for science,
time for ourselves and others. And, as things get the chance of going
slow and smooth, paradoxically, they go by without realising, and you
end up busy writing reports over a day you thought started instants
before. It’s a pleasure being so efficient and living so simply.

We’re only a third of the way through, but I’m already afraid it’ll be
over before I realise it. I’m enjoying every second of it while it lasts.

EVA Report – March 05th

Crew 223 EVA Report 05-03-2020

EVA # 4

Author: Marion Lebrun (Scientist)

Purpose of EVA: Change the batteries of MegaARES and LOAC. Explore a bit
further the surroundings of our Martian habitat and have a look at a
different kind of relief of the planet.

Start time: 8:50 AM

End time: 10:48 AM

Narrative: After the five minutes of depressurization, at 8:55 AM, we
walked out of the airlock. We checked the water level of the static tank
as well as the charge and usage duration of all 3 rovers. We wanted to
take the ATV Honda but we could not get it to start so we ended up
taking rover Spirit and ATVs 1 and 3.
We drove to our first stop which was right by the Hab where the LOAC and
MegaARES were set, to change the batteries of the two instruments. As we
were very organized, by 9:13 AM all the procedures were done and we
could leave for Candor Chasma.
We parked the vehicles before Steam Bed Connector and started the
exploration towards our final destination. The walk lasted about 35min.
It was quite surprising to see how sandy the area was. We also found
water on Mars during that EVA as well as antelope footsteps. The canyon
was truly beautiful and we were happy to find shade under some rocks
since it was getting really hot and sunny.
At 10:00 AM we decided to turn around to make sure we will on time at
the Hab. Unfortunately, on the way back, I lost my headset after
Aurélien tried to adjust my backpack that was hurting my back and my
neck. I had just enough time before I lost it to give the lead to Luc
who became the EVA Leader for the rest of the EVA.
We arrived at the place where we had left the vehicles at 10:31 AM and
we came back to the Hab at 10:43 AM to start the five minutes of
pressurization.

Destination: Plane area near the Hab along Cow Dung Road and the East
side of Candor Chasma

Coordinates (use UTM NAD27 CONUS): 12S 518500 4250800 and 12S 520750 4250889

Participants: Marion Lebrun (EVA Leader, Scientist), Florian Delpech
(Astronomer), Aurélien Mure (Commander), Luc Fortin (Engineer)

Road(s) and routes per MDRS Map: Cow Dung Road (driving) and Stream Bed
Connector (walking)

Mode of travel: Driving rover Spirit and ATVs 1 and 3 (and walking when
leaving the PEV road)